OpenStudy (anonymous):

what is 0.5, 0.519, 0.13, and 0.2 from smallest to largest

OpenStudy (anonymous):

.13, .2, .5, .519

OpenStudy (anonymous):

\[ \frac {130}{1000} <\frac {200}{1000} < \frac {500}{1000}< \frac {519}{1000} \]

OpenStudy (anonymous):

.13, .2, .5, .519 go by the number closet to the decimal point an treat it as the smallest

OpenStudy (anonymous):

thanks